OpenStudy (anonymous):

MEDAL AND FANNING WHO ANWSERS CORRECTLY FIRST!. Which answer shows this equation in standard form? 6 – 2(x – y) = –3x + 5 A. x + 2y = –1 B. x + 2y = 1 C. x – 2y = 1 D. –5x + 2y = –1

OpenStudy (anonymous):

b

OpenStudy (anonymous):

6-2(x-y)=-3x+5 6-2x+2y=-3x+5 6-2x+2x+2y=-3x+2x+5 6+2y=-x+5 6-6+2y=-x+5-6 2y=-x-1 2y+x=-x+x-1 x+2y=-1

OpenStudy (anonymous):

medal and fan plz :D

OpenStudy (anonymous):

@sydneymmelin ty and done :)

OpenStudy (bela):

You copied that from http://brainly.com/question/204733

OpenStudy (anonymous):

Like ur pic btw how do u get that

OpenStudy (anonymous):

@sydneymmelin lol thnx :)

OpenStudy (anonymous):

either way i answer correctly and u didnt answer at all @Bela she didnt say we couldnt find the answer

OpenStudy (anonymous):

duh

OpenStudy (anonymous):

@Bela I'm posting more questions so hopefuly I can fan and medal u as well :)

OpenStudy (anonymous):

@sydneymmelin lol

OpenStudy (bela):

I'm not stupid, I was trying to find the answer like a normal person instead of copying it so that's why you got the answer before me. Whatever.

OpenStudy (anonymous):

Lol sorry i am a normal person who just finds the answer its normal to look it up on google hes ur best friend